# Thermodynamics Reconstructed from Information Science: Theoretical Core

### v1

@@ Abstract
We present a minimal information-theoretic/axiomatic framework that reconstructs thermodynamic structure from a pair of primitive data: a measurement (coarse-graining) map $M$ and a reference measure (prior) $\tau$. For each cell $C_{y}=M^{-1}(y)$ in the state space, we introduce an information volume $\tau(C_{y})$ and its log-volume, and incorporate it—on the same footing as energy and particle number—as an expectation constraint in a minimum relative-entropy (Min-KL) problem. Temperature, chemical potential, and pressure are then defined primarily as dual variables of a relative-entropy functional, yielding a Legendre structure from which a first-law-type differential relation for internal energy follows. In particular, pressure is reinterpreted as the variable conjugate to an information volume rather than a geometric volume.
For nonequilibrium processes, we define the path-space relative entropy $\Sigma_{0,T}$ between forward and time-reversed path measures as entropy production at the path level, and decompose $\Sigma_{0,T}$ into the sum of the system entropy change and an environment entropy change using observational entropy $S_{M,\tau}$. With temperature introduced as the static Min-KL dual variable, we define heat as $Q:=T\,\Delta S_{\mathrm{env}}$, thereby reconstructing the conventional heat quantity as the representation $Q=T(\Sigma_{0,T}-\Delta S_{\mathrm{sys}})$ without model-dependent assumptions such as local detailed balance.
We further compare path-space relative entropies of a joint system and a subsystem, and show that their difference is expressed as an expectation of the KL divergence between conditional path measures, proving nonnegativity. This unifies Ito–Sagawa-type information terms and “hidden dissipation” as an information-geometric gap induced by projection from the joint description to a marginal one. On the static side, for additive Hamiltonians we formulate the exact decomposition of the joint nonequilibrium free energy into subsystem free energies plus a correlation free-energy term proportional to the mutual information $I(X;Y)$, understood as a KL-geometric decomposition.
Finally, we treat the choice of $(M,\tau)$ as a gauge (convention), distinguish gauge-invariant quantities (e.g. path relative entropy and free-energy differences) from gauge-dependent bookkeeping terms, and present a two-state Markov-chain toy model where $\Sigma_{0,T}$, observational dissipation, hidden dissipation, and heat $Q$ are computed explicitly.
@@ 1 Introduction
    We introduce an information-volume variable derived from primitive data $(M,\tau)$—a measurement map $M$ and a reference measure $\tau$—and define pressure as a primary dual variable, symmetric to temperature and chemical potential, through an expectation constraint on the log-volume.
@@
    We introduce dissipation as the path-space KL divergence $\Sigma_{0,T}=D_{\mathrm{KL}}\!\left(\mathbb{P}\,\middle\|\,\mathbb{P}^{\mathrm{R}}\right)$ in a primary manner, and define the environment entropy change $\Delta S_{\mathrm{env}}$ as a representation theorem when combined with observational entropy $S_{M,\tau}$.
@@
We present static Min-KL inference, dynamic path-KL dissipation, and the resulting information-theoretic representation of heat in as model-independent a form as possible.
2 Proposed model and theoretical framework: primitive data, information volume, and observational entropy
2.1 Primitive data
We fix the following three items as primitive data:
  1. 1.
    State space: a measurable space $(\Omega,\mathcal{F})$.
  2. 2.
    Measurement (coarse-graining): a measurable map $M:\Omega\to\mathcal{Y}$ (including finite resolution, partitions, and measurement).
  3. 3.
    Reference (coding convention / prior): a reference measure $\tau$ (assumed $\sigma$-finite).
The measurement map $M$ specifies which degrees of freedom are distinguished by observation. It associates a microstate $x\in\Omega$ with an observed outcome $y=M(x)$; thus $M$ represents, in a unified manner, “cell partitions of phase space,” “instrument resolution,” and “the choice of coarse-graining.” For example, if in a one-dimensional particle system the position is only measured on a mesh of size $\Delta x$, then $M$ may be regarded as mapping a continuous coordinate $x$ to a discrete label $y\in\mathbb{Z}$.
The reference measure $\tau$ encodes a prior (coding convention): “which states are regarded as uniform” at the descriptive level. For continuous systems, the Liouville or Lebesgue measure is a natural candidate; for discrete systems one often chooses the uniform distribution or an equilibrium distribution. In this paper, we introduce $\tau$ primarily as an information-theoretic reference measure, and assign physical meaning (equilibrium vs. nonequilibrium, etc.) as needed later.
Let the cell (equivalence class) induced by $M$ be
$C_{y}\;:=\;M^{-1}(y).$ (1)
The prior $\tau$ then induces a “volume” of each cell,
$v(y)\;:=\;\tau(C_{y}),$ (2)
which is not a geometric volume but a measure determined by the observation and coding convention. We call $v(y)$ the information volume.
Moreover, for each microstate $x$ we define the associated log-volume
@@ 2.1 Primitive data
This is the “information-volume” variable used throughout the paper.
@@ 2.2 Observational entropy
Given a microdistribution $p$ (with $p\ll\tau$), define the macrodistribution
$p_{M}(y)\;:=\;p(C_{y}).$ (4)
A representative form of observational entropy is
@@
for finite $\mathcal{Y}$. The framework and properties of observational entropy, including extensions to quantum settings and general priors, have been systematically developed by Safránek and collaborators [3, 4, 5, 6, 7]. In this paper, we take $S_{M,\tau}$ as the primary definition of “system entropy” in both static and dynamic discussions.
2.3 Gauge freedom in the choice of $M$ and $\tau$
The choice of the measurement map $M$ and the reference measure $\tau$ essentially fixes a gauge of description (a convention). We summarize which quantities depend on this choice and which are invariant.
(i) Refinement/coarse-graining of the measurement map
We say that $M_{1}:\Omega\to\mathcal{Y}_{1}$ is a refinement of $M_{2}:\Omega\to\mathcal{Y}_{2}$ if each equivalence class $C_{y}^{(2)}:=M_{2}^{-1}(y)$ is further partitioned into cells $C_{y^{\prime}}^{(1)}:=M_{1}^{-1}(y^{\prime})$. Then $M_{1}$ corresponds to a finer observation than $M_{2}$.
According to Safránek et al. [5], under suitable conditions one has
@@ 2.3 Gauge freedom in the choice of $M$ and $\tau$
Thus observational entropy can increase under coarse observation. Consequently, the system entropy change
@@
depends on the choice of $M$. Likewise, the environment entropy change $\Delta S_{\mathrm{env}}$ introduced in Sec. 5 also changes with $M$.
In contrast, the path-KL dissipation
$\Sigma_{0,T}=D_{\mathrm{KL}}\!\left(\mathbb{P}\,\middle\|\,\mathbb{P}^{\mathrm{R}}\right)$
is determined solely by the forward and reverse path measures $\mathbb{P},\mathbb{P}^{\mathrm{R}}$ and does not depend on $M$. In this sense, $\Sigma_{0,T}$ is gauge-invariant under changes of observation, while $\Delta S_{\mathrm{sys}}$ and $\Delta S_{\mathrm{env}}$ are gauge-dependent bookkeeping terms that allocate dissipation between “system” and “environment.”
(ii) Changing the reference measure and invariant content
Next, consider changing the reference measure from $\tau$ to $\tau^{\prime}$. Assume $\tau$ and $\tau^{\prime}$ are mutually absolutely continuous, so that for some integrable $\varphi$,
$\mathrm{d}\tau^{\prime}(x)=e^{-\varphi(x)}\,\,\mathrm{d}\tau(x).$
Then for any distribution $p$,
$D_{\mathrm{KL}}\!\left(p\,\middle\|\,\tau^{\prime}\right)=D_{\mathrm{KL}}\!\left(p\,\middle\|\,\tau\right)+\int\varphi(x)\,p(\,\mathrm{d}x)+\text{(constant)}.$
Hence the Min-KL structure is preserved, while the numerical value of the optimum $D_{\mathrm{KL}}\!\left(p^{\ast}\,\middle\|\,\tau\right)$ generally depends on $\tau$.
In this paper we fix $\tau$ as a coding convention specifying “uniformity” and define $S_{M,\tau}$, $\ell_{V}$, pressure $P$, etc. consistently under that choice. Switching to another $\tau^{\prime}$ corresponds to a gauge change of the theory; physical content can be extracted by focusing on quantities invariant under this change (e.g. entropy differences between two states, and the path-KL dissipation $\Sigma_{0,T}$).
@@
both $\Delta S_{\mathrm{sys}}$ and $\Delta S_{\mathrm{env}}$ depend on the choice of $M$ and $\tau$. However, $\Sigma_{0,T}$ itself is uniquely determined once the forward and reverse path measures are given and is invariant under gauge changes of $(M,\tau)$. Thus our framework separates the total dissipation from its allocation between system and environment.
@@ 3.1 Variational problem (state as a definition)
$\mathbb{E}_{p}[E]=U,\qquad\mathbb{E}_{p}[N]=\bar{N},\qquad\mathbb{E}_{p}[\ell_{V}]=\bar{v}.$ (6)
@@
$D_{\mathrm{KL}}\!\left(p\,\middle\|\,\tau\right)\;:=\;\int_{\Omega}\log r(x)\,p(\,\mathrm{d}x)\;=\;\int_{\Omega}r(x)\log r(x)\,\tau(\,\mathrm{d}x),\qquad r=\frac{\,\mathrm{d}p}{\,\mathrm{d}\tau}.$ (8)
3.2 Exponential-family solution and dual variables
Under standard convexity and regularity conditions, the solution belongs to an exponential family:
@@ 3.2 Exponential-family solution and dual variables
where $Z$ is the normalization constant. The multipliers $\beta,\alpha,\pi$ are Lagrange multipliers (shadow prices / dual variables); general conjugacy relations for continuous MaxEnt are discussed in [8].
@@ 3.3 Definitions of temperature, chemical potential, and pressure
Our policy is to define temperature, chemical potential, and pressure primarily as dual variables, rather than introducing them a posteriori (we set $k_{\mathrm{B}}=1$):
@@
$\mu$: shadow price of the particle-number constraint $\mathbb{E}[N]=\bar{N}$.
$P$: shadow price of the information-volume constraint $\mathbb{E}[\ell_{V}]=\bar{v}$.
In this sense, pressure $P$ is defined information-theoretically as the increase rate of KL with respect to the log-volume constraint, without relying on a geometric notion of volume.
3.4 Dual potential and a first-law-type relation
@@
$\Psi(\beta,\alpha,\pi)\;:=\;\log\int_{\Omega}\exp\!\bigl(-\beta E(x)-\alpha N(x)-\pi\ell_{V}(x)\bigr)\,\tau(\,\mathrm{d}x)\;=\;\log Z(\beta,\alpha,\pi).$ (11)
Under regularity,
$U=-\partial_{\beta}\Psi,\qquad\bar{N}=-\partial_{\alpha}\Psi,\qquad\bar{v}=-\partial_{\pi}\Psi.$ (12)
Let the optimal value (maximum relative entropy) be
$S^{\ast}(U,\bar{N},\bar{v})\;:=\;-D_{\mathrm{KL}}\!\left(p^{\ast}\,\middle\|\,\tau\right).$ (13)
Convex duality yields the differential relation
$\mathrm{d}S^{\ast}=\beta\,\,\mathrm{d}U+\alpha\,\,\mathrm{d}\bar{N}+\pi\,\,\mathrm{d}\bar{v}.$ (14)
Substituting (10) gives
$\mathrm{d}U=T\,\,\mathrm{d}S^{\ast}-P\,\,\mathrm{d}\bar{v}+\mu\,\,\mathrm{d}\bar{N}.$ (15)
@@
Assume the Min-KL problem (7) has a unique solution $p^{\ast}$, and the corresponding log-partition function $\Psi(\beta,\alpha,\pi)$ is a $C^{2}$ convex function. Let the optimal value $S^{\ast}(U,\bar{N},\bar{v})$ be defined by (13). Then the total differential in $(U,\bar{N},\bar{v})$ satisfies (14). Moreover, if intensive variables are defined by (10), the total differential of internal energy $U$ satisfies (15).
@@
By convex duality, $S^{\ast}$ is given by the Legendre transform
$S^{\ast}(U,\bar{N},\bar{v})=\inf_{\beta,\alpha,\pi}\bigl\{\beta U+\alpha\bar{N}+\pi\bar{v}-\Psi(\beta,\alpha,\pi)\bigr\}.$
Under regularity, the optimal $(\beta,\alpha,\pi)$ is unique, and equivalently to (12) one has $\partial_{U}S^{\ast}=\beta$, $\partial_{\bar{N}}S^{\ast}=\alpha$, $\partial_{\bar{v}}S^{\ast}=\pi$. Hence the total differential is (14).
Substituting (10) yields
$\mathrm{d}S^{\ast}=\frac{1}{T}\,\,\mathrm{d}U-\frac{\mu}{T}\,\,\mathrm{d}\bar{N}+\frac{P}{T}\,\,\mathrm{d}\bar{v}.$
Multiplying by $T$ and solving for $\mathrm{d}U$ gives (15). ∎
@@ 4 Nonequilibrium extension: path-space KL and hidden dissipation
We extend the preceding information-theoretic framework to nonequilibrium processes with time evolution. Irreversibility (entropy production, dissipation) is naturally defined as time-reversal asymmetry of stochastic processes; the most general coordinate-free formulation is given by the KL divergence between the forward and time-reversed path measures [9, 10, 11].
@@ 4.1 Definition of path dissipation (entropy production)
Consider a stochastic process defined on a time interval $[0,T]$. Let $\mathbb{P}$ be the forward path measure on path space, and let $\mathbb{P}^{\mathrm{R}}$ be the path measure of the time-reversed process induced by the time-reversal map. We define
$\Sigma_{0,T}\;:=\;D_{\mathrm{KL}}\!\left(\mathbb{P}\,\middle\|\,\mathbb{P}^{\mathrm{R}}\right)\;\geq 0$ (16)
as entropy production (dissipation) on $[0,T]$. This definition does not invoke any model-specific assumptions such as bath temperature or LDB.
@@ 4.2 Lower bound under observation (coarse-graining) and hidden dissipation
We represent observation as a path map (channel)
$\Pi:\ \omega=(x_{0:T})\longmapsto\gamma=(y_{0:T}),\qquad y_{t}=M(x_{t}),$ (17)
where $M$ is the measurement map in (1). Write the observed path measures as $\Pi\mathbb{P}$ and $\Pi\mathbb{P}^{\mathrm{R}}$. Define observational dissipation (entropy production at the observation level) by
@@
Then by the data-processing inequality,
$\Sigma_{0,T}=D_{\mathrm{KL}}\!\left(\mathbb{P}\,\middle\|\,\mathbb{P}^{\mathrm{R}}\right)\;\geq\;D_{\mathrm{KL}}\!\left(\Pi\mathbb{P}\,\middle\|\,\Pi\mathbb{P}^{\mathrm{R}}\right)=\Sigma_{\mathrm{obs}}$ (19)
holds [12, 13].
Next consider a joint system $(X,Y)$. Let $\mathbb{P}_{XY},\mathbb{P}_{XY}^{\mathrm{R}}$ be the forward and reverse path measures of the joint system, and let $\mathbb{P}_{X},\mathbb{P}_{X}^{\mathrm{R}}$ be those of the subsystem $X$. Define the joint and marginal dissipations by
$\Sigma_{XY}$ $:=D_{\mathrm{KL}}\!\left(\mathbb{P}_{XY}\,\middle\|\,\mathbb{P}_{XY}^{\mathrm{R}}\right),$ (20)
$\Sigma_{X}$ $:=D_{\mathrm{KL}}\!\left(\mathbb{P}_{X}\,\middle\|\,\mathbb{P}_{X}^{\mathrm{R}}\right).$ (21)
Their difference
@@
is called the “hidden dissipation from the viewpoint of XX.”
@@
Assume the joint forward and reverse path measures $\mathbb{P}_{XY},\mathbb{P}_{XY}^{\mathrm{R}}$ admit densities with respect to a common reference measure, and conditional path measures $\mathbb{P}_{Y|X},\mathbb{P}_{Y|X}^{\mathrm{R}}$ exist. Then the hidden dissipation $\Sigma_{\mathrm{hidden}}^{(X)}$ admits the representation
$\Sigma_{\mathrm{hidden}}^{(X)}=\int D_{\mathrm{KL}}\!\left(\mathbb{P}_{Y|X=\gamma}\,\middle\|\,\mathbb{P}_{Y|X=\gamma}^{\mathrm{R}}\right)\,\mathbb{P}_{X}(\,\mathrm{d}\gamma)\;\geq\;0.$ (23)
That is, the difference between joint and marginal dissipation is the expectation of the KL divergence between conditional path measures, and is always nonnegative.
@@
Write the joint path measures as $\mathrm{d}\mathbb{P}_{XY}=\,\mathrm{d}\mathbb{P}_{X}\,\,\mathrm{d}\mathbb{P}_{Y|X}$ and $\mathrm{d}\mathbb{P}_{XY}^{\mathrm{R}}=\,\mathrm{d}\mathbb{P}_{X}^{\mathrm{R}}\,\,\mathrm{d}\mathbb{P}_{Y|X}^{\mathrm{R}}$. By the chain rule for KL divergence,
$\Sigma_{XY}$ $=\int\log\!\left(\frac{\,\mathrm{d}\mathbb{P}_{X}}{\,\mathrm{d}\mathbb{P}_{X}^{\mathrm{R}}}\right)\,\mathrm{d}\mathbb{P}_{X}+\int\left[\int\log\!\left(\frac{\,\mathrm{d}\mathbb{P}_{Y|X}}{\,\mathrm{d}\mathbb{P}_{Y|X}^{\mathrm{R}}}\right)\,\mathrm{d}\mathbb{P}_{Y|X}\right]\,\mathrm{d}\mathbb{P}_{X}$
$=\Sigma_{X}+\int D_{\mathrm{KL}}\!\left(\mathbb{P}_{Y|X=\gamma}\,\middle\|\,\mathbb{P}_{Y|X=\gamma}^{\mathrm{R}}\right)\,\mathbb{P}_{X}(\,\mathrm{d}\gamma).$ (24)
The integrand is a KL divergence and hence nonnegative; integrating it against $\mathbb{P}_{X}$ preserves nonnegativity, proving (23). ∎
This proposition clarifies that information terms and mutual-information flow appearing in Ito–Sagawa and Horowitz–Esposito-type frameworks can be interpreted as concrete closed-form evaluations of $\Sigma_{\mathrm{hidden}}^{(X)}$ in specific model classes [14, 15, 16].

### v2

# Thermodynamics Reconstructed from Information Theory: An Axiomatic Framework via Information-Volume Constraints and Path-Space KL Divergence

@@ Abstract
We develop an axiomatic reconstruction of thermodynamics based entirely on two primitive components: a description of what aspects of a system are observed and a reference measure that encodes the underlying descriptive convention. These ingredients define an “information volume” for each observational cell. By incorporating the logarithm of this volume as an additional constraint in a minimum–relative-entropy inference scheme, temperature, chemical potential, and pressure arise as conjugate variables of a single information-theoretic functional. This leads to a Legendre-type structure and a first-law-like relation in which pressure corresponds to information volume rather than geometric volume.
For nonequilibrium dynamics, entropy production is characterized through the relative-entropy asymmetry between forward and time-reversed stochastic evolutions. A decomposition using observational entropy then separates total dissipation into system and environment contributions. Heat is defined as the part of dissipation not accounted for by the system-entropy change, yielding a representation that does not rely on local detailed balance or a specific bath model. We further show that the difference between joint and partially observed dissipation equals the average of conditional relative entropies, providing a unified interpretation of hidden dissipation and information-flow terms as projection-induced gaps.
Keywords: information thermodynamics, KL divergence, observational entropy, information volume, dissipation, coarse-graining
@@ 1 Introduction
    We introduce an information–volume variable derived from the primitive data $(M,\tau)$—a measurement map $M$ and a reference measure $\tau$—and define pressure as a primary dual variable, on par with temperature and chemical potential, via an expectation constraint on the log-volume.
@@
    We introduce dissipation in a primary manner as the path-space KL divergence between the forward path measure $\mathbb{P}$ and its time-reversed counterpart $\mathbb{P}^{\mathrm{R}}$,
    $\Sigma_{0,T}:=D_{\mathrm{KL}}\!\left(\mathbb{P}\,\middle\|\,\mathbb{P}^{\mathrm{R}}\right)\,.$
    Combining this with the observational entropy $S_{M,\tau}$, we obtain a representation theorem that identifies the environmental entropy change $\Delta S_{\mathrm{env}}$.
@@
First, we start only from the primitive descriptive data $(M,\tau)$ and KL-based inference principles, we derive (i) an equilibrium state family as a Min-KL solution, (ii) a Legendre-type dual structure, (iii) the intensive variables $(T,\mu,P)$ as Lagrange multipliers conjugate to $(U,\bar{N},\bar{\ell}_{V})$, and (iv) a model-independent representation of heat from the pair (path-KL dissipation, observational entropy). Second, we identify these derived quantities with conventional thermodynamic nomenclature in regimes where additional structure is available (e.g. geometric volume or local detailed balance), while outside such regimes the same definitions remain internally consistent and comparable across descriptions. Consequently, because both static and dynamic notions are anchored to $(M,\tau)$ and path measures, the framework (a) defines pressure and heat even when geometric volume or a heat bath model are unavailable, (b) makes the system–environment entropy split explicitly gage dependent while keeping total dissipation gage invariant, (c) unifies hidden dissipation and information-flow terms as KL projection gaps, and (d) turns “observation design” into a principled problem (choosing $(M,\tau)$ to tighten observable dissipation bounds). As follows, we present static Min-KL inference, dynamic path-KL dissipation, and the resulting information-theoretic representation of heat in as model-independent a form as possible.
2 Primitive data, information volume, and observational entropy
We fix: (1) a measurable state space $(\Omega,\mathcal{F})$; (2) a measurable map $M:\Omega\to\mathcal{Y}$ specifying observation/coarse-graining; (3) a $\sigma$-finite reference measure $\tau$.
Let $C_{y}:=M^{-1}(y)$ be the induced cells. Their $\tau$-measures
$v(y):=\tau(C_{y})$ (1)
define an information volume. The associated log-volume variable is
@@ 2.1 Observational entropy
For $p\ll\tau$, define $p_{M}(y):=p(C_{y})$. Observational entropy is
@@
for finite $\mathcal{Y}$. The framework and properties of observational entropy—including quantum generalizations and multi-resolution extensions—have been developed in [5, 6, 7, 8, 9]. We adopt $S_{M,\tau}$ as the primary definition of system entropy.
2.2 Gauge freedom in the choice of $(M,\tau)$
Changes in $(M,\tau)$ represent changes of description.
2.2.1 (i) Refinement/coarse-graining of $M$.
If $M_{1}$ is a refinement of $M_{2}$ then (Safránek et al.)
@@ 2.2 Gauge freedom in the choice of $(M,\tau)$
Thus the system entropy change
@@
depends on $M$. However, the path-KL dissipation
$\Sigma_{0,T}=D_{\mathrm{KL}}(\mathbb{P}\|\mathbb{P}^{\mathrm{R}})$
is independent of $M$.
(ii) Changing the reference measure $\tau$.
If $\mathrm{d}\tau^{\prime}=e^{-\varphi}\,\mathrm{d}\tau$, then
$D_{\mathrm{KL}}\!\left((\,\middle\|\,p\right)\|\tau^{\prime})=D_{\mathrm{KL}}\!\left((\,\middle\|\,p\right)\|\tau)+\int\varphi\,\,\mathrm{d}p+\mathrm{const}.$
Thus changing $\tau$ corresponds to an additive gauge transformation of entropy.
@@
both $\Delta S_{\mathrm{sys}}$ and $\Delta S_{\mathrm{env}}$ depend on $(M,\tau)$, but $\Sigma_{0,T}$ is invariant. Our framework separates the gauge-invariant dissipation from the gauge-dependent bookkeeping of system/environment entropy.
@@ 3 Static theory: Min-KL inference and dual variables
$\mathbb{E}_{p}[E]=U,\qquad\mathbb{E}_{p}[N]=\bar{N},\qquad\mathbb{E}_{p}[\ell_{V}]=\bar{\ell}_{V}.$ (4)
@@
$D_{\mathrm{KL}}\!\left(p\,\middle\|\,\tau\right):=\int_{\Omega}\log r(x)\,p(\,\mathrm{d}x)=\int_{\Omega}r(x)\log r(x)\,\tau(\,\mathrm{d}x).$ (6)
Under convexity and regularity assumptions, the minimizer belongs to the exponential family:
@@
where $Z$ is the normalization constant.
@@ 3.1 Definitions of temperature, chemical potential, and pressure
We define the intensive variables primarily as dual variables:
@@
Thus temperature is the multiplier of the energy constraint, chemical potential is associated with the particle-number constraint, and pressure is the shadow price of the information-volume constraint.
Intuition and a non-geometric example.
The “information volume” $v(y)=\tau(C_{y})$ is the size (measure) of the set of microstates compatible with a given observational outcome $y$. Unlike geometric volume, $v(y)$ can quantify degeneracy or description-dependent resolution in state spaces that are discrete, combinatorial, or otherwise non-spatial. Constraining $\mathbb{E}[\ell_{V}]$ with $\ell_{V}=\log v(M(x))$ fixes the typical code length associated with the observational cells visited by the system, and the multiplier $\pi=P/T$ is the corresponding shadow price: it quantifies how strongly the Min-KL state penalizes occupying outcomes with large degeneracy (large $v$), even when no physical container volume exists.
Example (combinatorial degeneracy).
Let $\Omega=\{0,1\}^{N}$ (spin configurations) with counting measure $\tau$, and let $M$ output the magnetization $m=\frac{1}{N}\sum_{i}(2x_{i}-1)$. Then each cell $C_{m}$ consists of all configurations with magnetization $m$, hence $v(m)=|C_{m}|=\binom{N}{(N(1+m))/2}$ and $\ell_{V}=\log v(m)$ is purely combinatorial. The constraint $\mathbb{E}[\ell_{V}]=\bar{\ell}_{V}$ controls the typical degeneracy of the observed macrostates, and $P$ is the intensive variable conjugate to this non-geometric “volume” in the same dual sense as $T$ and $\mu$.
@@
$\Psi(\beta,\alpha,\pi):=\log\int_{\Omega}\exp\!\Bigl(-\beta E-\alpha N-\pi\ell_{V}\Bigr)\,\tau(\,\mathrm{d}x)=\log Z(\beta,\alpha,\pi).$ (9)
Then
$U=-\partial_{\beta}\Psi,\qquad\bar{N}=-\partial_{\alpha}\Psi,\qquad\bar{\ell}_{V}=-\partial_{\pi}\Psi.$ (10)
Let
$S^{\ast}(U,\bar{N},\bar{\ell}_{V}):=-D_{\mathrm{KL}}\!\left(p^{\ast}\,\middle\|\,\tau\right).$ (11)
Convex duality gives
$\mathrm{d}S^{\ast}=\beta\,\,\mathrm{d}U+\alpha\,\,\mathrm{d}\bar{N}+\pi\,\,\mathrm{d}\bar{\ell}_{V}.$ (12)
Substituting (8) yields a first-law-type differential relation:
$\mathrm{d}U=T\,\,\mathrm{d}S^{\ast}-P\,\,\mathrm{d}\bar{\ell}_{V}+\mu\,\,\mathrm{d}\bar{N}.$ (13)
@@
Assume the Min-KL problem (5) has a unique solution $p^{\ast}$, and that $\Psi(\beta,\alpha,\pi)$ is $C^{2}$ and convex. Let $S^{\ast}$ be defined by (11). Then the total differential satisfies (12). With the intensive variables defined by (8), the differential of internal energy satisfies (13).
@@
By convex duality,
$S^{\ast}(U,\bar{N},\bar{\ell}_{V})=\inf_{\beta,\alpha,\pi}\{\beta U+\alpha\bar{N}+\pi\bar{\ell}_{V}-\Psi(\beta,\alpha,\pi)\}.$
The minimizer satisfies the stationarity conditions (10), giving
$\partial_{U}S^{\ast}=\beta,\qquad\partial_{\bar{N}}S^{\ast}=\alpha,\qquad\partial_{\bar{\ell}_{V}}S^{\ast}=\pi,$
hence (12). Inserting (8) and solving for $\mathrm{d}U$ yields (13). ∎
@@ 4 Nonequilibrium extension: path-space KL and hidden dissipation
We extend the preceding information-theoretic framework to time-evolving nonequilibrium processes. Irreversibility (entropy production, dissipation) is defined as the time-reversal asymmetry of path measures. The most general formulation—independent of coordinate choices and independent of any model-specific assumptions such as local detailed balance—is the path-space KL divergence [1, 2, 3].
@@ 4.1 Definition of path dissipation (entropy production)
Let $\mathbb{P}$ be the forward path measure of a stochastic process on $[0,T]$, and $\mathbb{P}^{\mathrm{R}}$ the time-reversed path measure induced by the reversal map. Define dissipation by
This definition requires no thermodynamic model, no heat bath, and no LDB assumptions.
@@ 4.2 Lower bound under observation (coarse-graining) and hidden dissipation
Observation in time induces a map on paths:
$\Pi:\ (x_{0:T})\mapsto(y_{0:T}),\qquad y_{t}=M(x_{t}).$ (15)
Define observed path measures $\Pi\mathbb{P}$ and $\Pi\mathbb{P}^{\mathrm{R}}$, and observed dissipation
@@
By the data-processing inequality,
$\Sigma_{0,T}=D_{\mathrm{KL}}\!\left(\mathbb{P}\,\middle\|\,\mathbb{P}^{\mathrm{R}}\right)\geq D_{\mathrm{KL}}\!\left(\Pi\mathbb{P}\,\middle\|\,\Pi\mathbb{P}^{\mathrm{R}}\right)=\Sigma_{\mathrm{obs}}.$ (17)
4.2.1 Joint / marginal dissipations and hidden dissipation.
Consider a joint system $(X,Y)$ with joint forward and reverse path measures $\mathbb{P}_{XY},\mathbb{P}_{XY}^{\mathrm{R}}$, and marginal measures $\mathbb{P}_{X},\mathbb{P}_{X}^{\mathrm{R}}$. Define
$\Sigma_{XY}$ $:=D_{\mathrm{KL}}(\mathbb{P}_{XY}\|\mathbb{P}_{XY}^{\mathrm{R}}),$ (18)
$\Sigma_{X}$ $:=D_{\mathrm{KL}}(\mathbb{P}_{X}\|\mathbb{P}_{X}^{\mathrm{R}}).$ (19)
The difference
@@
quantifies the dissipation invisible from subsystem $X$.
@@
Assume that joint and marginal path measures admit densities with respect to common references, and that conditional path measures $\mathbb{P}_{Y|X},\mathbb{P}^{\mathrm{R}}_{Y|X}$ exist. Then
$\Sigma_{\mathrm{hidden}}^{(X)}=\int D_{\mathrm{KL}}\!\left(\mathbb{P}_{Y|X=\gamma}\,\middle\|\,\mathbb{P}^{\mathrm{R}}_{Y|X=\gamma}\right)\mathbb{P}_{X}(\,\mathrm{d}\gamma)\geq 0.$ (21)
@@
Write
$\mathrm{d}\mathbb{P}_{XY}=\,\mathrm{d}\mathbb{P}_{X}\,\,\mathrm{d}\mathbb{P}_{Y|X},\qquad\,\mathrm{d}\mathbb{P}^{\mathrm{R}}_{XY}=\,\mathrm{d}\mathbb{P}^{\mathrm{R}}_{X}\,\,\mathrm{d}\mathbb{P}^{\mathrm{R}}_{Y|X}.$
Then apply the chain rule for KL divergence:
$\Sigma_{XY}$ $=\int\log\!\frac{\,\mathrm{d}\mathbb{P}_{X}}{\,\mathrm{d}\mathbb{P}^{\mathrm{R}}_{X}}\,\mathrm{d}\mathbb{P}_{X}+\int\left[\int\log\!\frac{\,\mathrm{d}\mathbb{P}_{Y|X}}{\,\mathrm{d}\mathbb{P}^{\mathrm{R}}_{Y|X}}\,\mathrm{d}\mathbb{P}_{Y|X}\right]\,\mathrm{d}\mathbb{P}_{X}$
$=\Sigma_{X}+\int D_{\mathrm{KL}}(\mathbb{P}_{Y|X=\gamma}\|\mathbb{P}^{\mathrm{R}}_{Y|X=\gamma})\mathbb{P}_{X}(\,\mathrm{d}\gamma).$ (22)
The KL divergence is nonnegative, completing the proof. ∎
This result shows that “hidden dissipation” is not an ad hoc correction but a KL projection gap. It provides a geometric foundation for mutual-information flow terms in frameworks of Ito–Sagawa, Horowitz–Esposito, etc. [14, 15, 16].